Grade 11Discuss with colleagues and IITians

I wish to have more interactive session of Mole Concepts - Molarity, Molality, Mole Fraction and Limiting reagents and the relation between them. I could not understand clearly from the study material available. I want to solve a problem also in this.

Question:

Calculate the Molarity of a solution of Ethanol in water, in which the mole fraction of Etanol is 0.040?

Please send me the solution

Gouthuam Lakshminarayanan
14 Years agoGrade 11

2 Answers

JAGANNATH SAHOO
14 Years ago

molarity M= (WB/WA)*1000/V

mole fraction of C2H5OH= 0.040

molar wt. of ethanol= 46gm

wt. of ethanol in 1 mole of soln= 46*0.04=1.84

mole fraction of water= 0.960

molar wt. of water= 18gm

wt. of water in 1 mole of solution=17.28

Volume of solution= 1.84+17.28= 19.12ml (1gm=1ml)

therefore,

M= (1.84/17.28)*1000/0.01921

=0.00556 M
thank you

Gouthuam Lakshminarayanan
14 Years ago

it should be 19.12 not 0.01912 as we are taking it in ml not in litres therefore on calculation answer should 

M= (1.84/17.28)*1000/19,12 = 5.56 M 

hope i am correct